OpenStudy (anonymous):

1. Which of the following is the value of x in the solution to the system of equations given below? 8 + 2x =5y 4x - y = 2 A)1 B)2 C)3 D)4

OpenStudy (anonymous):

Multiply the second equation with " -5 " and add the two equations. In this way you can eliminate y and solve for x

OpenStudy (mathstudent55):

Do what @bhaskarbabu wrote above, but first, subtract 2 and add y to both sides of the second equation, so you have the two equations properly lined up.

OpenStudy (anonymous):

No just multiply the entire second equation with " -5"

OpenStudy (mathstudent55):

8 + 2x =5y ---> 8 + 2x = 5y 4x - y = 2 ---> -2 + 4x = y Now that the like terms are lined up properly, multiply the second equation by -5 and add them together. 8 + 2x = 5y ---> 8 + 2x = 5y -2 + 4x = y ---> 10 - 20x = -5y -------------- 18 - 18x = 0 Now solve for x.